thank you thumbs up!





3. Solve 2x2y – 9xy +12y = 0.
Answer #1

Given differendial eqr is 2x²y - guyti2y=0 - 0 Z put 1=e - Inu=2 € dz an 1 2 dz mou dy dy dz ZP v are np LP IP KE xp - dz nsolution of of ei is 42 CZ ci e 412 e Y12) z=Inse by back substitution 4lok => Yl2)=ci e (32) Inne + c2 e 312 2 + C2 => Y (21